# On Focusing Statistical Power for Searches and Measurements in Particle Physics

### v1

@@ Acknowledgements.
This project was initiated through discussions at the PHYSTAT workshops in Munich and London in 2024, and we thank Mikael Kuusela, David Rousseau, Ibles Olcina, Callum Wilkinson, and Lukas Heinrich for fruitful conversations. We thank Scott Kravitz for guidance on setting up the LZ-inspired dataset, pointing us to relevant resources, and for feedback on the manuscript. AG and DW were supported by the DOE Office of Science. JC and ABL were supported in part by NSF DMS-2053804. RI is grateful for the financial support of FAPESP (grant 2023/07068-1) and CNPq (grants 422705/2021-7 and 305065/2023-8). This work used Bridges-2 at PSC through allocation MTH250002 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program [33], which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ Appendix A Experimental setup
The above “counting experiment” model convention is widely used for making likelihood-free inference tractable in cases where rare or exotic events are mixed with a relatively high volume of background events [34, 35]. In this Letter, we follow the standard convention of histogram analyses for such counting experiments.
@@ A.1 Implementation of test statistics
This adjustment mitigates the effect of the boundary on the distribution of the LRS itself, independent of the method used for critical value estimation. The boundary effect arises because Wilks’ theorem, which guarantees that the critical values remain constant across the parameter space for LRS and large sample sizes (as in Figure 1 top), only holds for interior points [36, 37]. To ensure that the likelihood is well defined on $\mu<0$, we amend Eq. 4 to express
$p(\mathcal{D}|\mu,{\bm{\alpha}},{\bm{\beta}})=\prod_{i=1}^{d}f(N_{i};\lvert\mu\alpha_{i}+\beta_{i}\rvert)$ (6)
with positive rate. The supremum is computed using scipy’s implementation of the bounded Brent routine [38].
@@ A.2 LZ-inspired dataset
We construct a dataset with signal and background shapes visually similar to those in Fig. 3 of Ref. [3]. Template histograms are generated using the nestpy Python interface (v2.0.4) to the NEST simulator (v2.4.0). The Lux_Run03 detector binding is adapted for this study, with the default parameter values used to determine the background electronic recoil (ER) template histogram and with set values of $g_{1}=0.092$ and $g_{1}^{gas}=0.076$ for the signal nuclear recoil (NR) histogram. These values were chosen to visually match the 40 GeV WIMP signal distribution in Ref. [3]. The resulting histograms are shown in Fig. 4. more detailed reproduction of the LZ setup or the use of official LZ simulation samples is beyond the scope of this Letter.
@@ Appendix B Quantile regression
To perform quantile regression, we generate pseudo-experiments to create a calibration dataset $\mathcal{T}=\{(\mu_{1},\mathcal{D}_{1}),\ldots,(\mu_{B},\mathcal{D}_{B})\}$ according to Eq. 4 from $\mu_{i}\sim\text{Unif}(\widetilde{\Theta})$, where $\widetilde{\Theta}$ is given by Eq. 5. For a code implementation of the Neyman construction using probabilistic regression along with examples for a range of test statistics, refer to https://github.com/lee-group-cmu/lf2i [12].

### v2

@@ Acknowledgements.
This project was initiated through discussions at the PHYSTAT workshops in Munich and London in 2024, and we thank Mikael Kuusela, David Rousseau, Ibles Olcina, Callum Wilkinson, and Lukas Heinrich for fruitful conversations. We thank Scott Kravitz for guidance on setting up the LZ-inspired dataset and pointing us to relevant resources. We also thank Scott Kravitz and Ben Nachman for feedback on the manuscript. AG and DW were supported by the DOE Office of Science. JC and ABL were supported in part by NSF DMS-2053804. RI is grateful for the financial support of FAPESP (grant 2023/07068-1) and CNPq (grants 422705/2021-7 and 305065/2023-8). This work used Bridges-2 at PSC through allocation MTH250002 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program [10.1145/3569951.3597559], which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ Appendix A Experimental setup
The above “counting experiment” model convention is widely used for making likelihood-free inference tractable in cases where rare or exotic events are mixed with a relatively high volume of background events [cowan1998statistical, casadei_reference_2012]. In this Letter, we follow the standard convention of histogram analyses for such counting experiments.
@@ A.1 Implementation of test statistics
This adjustment mitigates the effect of the boundary on the distribution of the LRS itself, independent of the method used for critical value estimation. The boundary effect arises because Wilks’ theorem, which guarantees that the critical values remain constant across the parameter space for LRS and large sample sizes (as in Figure 1 top), only holds for interior points [Algeri:2019lah, Herold:2024enb]. To ensure that the likelihood is well defined on $\mu<0$, we amend Eq. 4 to express
[…]
with rate parameter ensured to be positive. The supremum is computed using scipy’s implementation of the bounded Brent routine [2020SciPy-NMeth].
@@ A.2 LZ-inspired dataset
We construct a dataset with signal and background shapes visually similar to those in Fig. 3 of Ref. [LZ:2024zvo]. Template histograms are generated using the nestpy Python interface (v2.0.4) to the NEST simulator (v2.4.0). The Lux_Run03 detector binding is adapted for this study, with the default parameter values used to determine the background electronic recoil (ER) template histogram and with set values of $g_{1}=0.092$ and $g_{1}^{gas}=0.076$ for the signal nuclear recoil (NR) histogram. These values were chosen to visually match the 40 GeV WIMP signal distribution in Ref. [LZ:2024zvo]. The resulting histograms are shown in Fig. 4. more detailed reproduction of the LZ setup or the use of official LZ simulation samples is beyond the scope of this Letter.
@@ Appendix B Quantile regression
[…]